# GenMix: Effective Data Augmentation with Generative Diffusion Model Image Editing

### v3

@@ 5.6 Domain Adaptation
When DNNs encounter data from distributions that differ from their training data, a noticeable decline in performance is often observed. To tackle this issue, domain adaptation framework Ganin et al. (2016); Kouw and Loog (2019) has been developed, which utilizes labeled data from source domains to train DNNs for unlabeled target domains using a transductive learning approach. We validate the effectiveness of GMix using LEAD framework Qu et al. (2024) on OfficeHome Venkateswara et al. (2017) dataset. For a fair comparison, we access source-only and augment data with GMix. We experiment with different settings including across partial-set domain adaptation (PDA), open-set domain adaptation (OSDA), and open-partial-set domain adaptation (OPDA). The OPDA results are reported in Table VI, while OSDA and PDA performance evaluations are summarized in Tables VII and VIII, respectively.
@@
Open-Partial Domain Adaptation (OPDA). OPDA Saito and Saenko (2021b); You et al. (2019) falls under the broader scope of universal domain adaptation, where there is no prior knowledge about the label shift, such as common classes or the number of categories in the target domain. Some existing approaches Saito et al. (2020b); Chen et al. (2022c) require simultaneous access to both source and target data which may become impractical due to data protection regulations Voigt and Von dem Bussche (2017). In Table VI, our baseline LEAD Qu et al. (2024) has an average H-score of $75.0\%$, while LEAD + GMix achieves $77.2\%$, representing a significant $2.2\%$ improvement in the average H-score. Particularly, for Ar2Pr, addition of our augmentation strategy improved performance of LEAD Qu et al. (2024) by $4.9\%$, for CL2Ar the improvement is $6.2\%$ and for Re2Ar improvement is $4.4\%$. The results affirm that GMix is an effective data augmentation scheme for open partial domain adaptation scenarios.
@@
Open-set Domain Adaptation (OSDA). In OSDA a significant portion of the target samples comes from new categories that are absent in the source domain. Since these categories do not have specific labels, existing methods often treat them collectively as a single ”unknown” class. In Table VII, LEAD + GMix achieved an average improvement in the H-score of $1.5\%$ compared to the baseline LEAD, underscoring the significance of our proposed augmentation strategy. Particularly, considerable improvements of $4.1\%$ for Cl2Ar, $5.4\%$ for Re2CL and $3.4\%$ for Re2Pr; can be observed by the addition of GMix augmentation in the baseline LEAD Qu et al. (2024) method.
@@
Partial Domain Adaptation (PDA). PDA focuses on adapting a model trained on a source domain to a target domain where only a subset of the classes overlap. We compare the performance of a number of techniques on the Office-Home dataset while adapting LEAD as LEAD + GMixto showcase the gain enabled by our method. Table VIII shows the accuracy for various domain adaptation tasks as well as the average accuracy across all tasks. LEAD + GMix achieved an average improvement in the H-score of $0.5\%$ over the LEAD. Particularly, for Pr2Cl maximum improvement of $9.5\%$, in Re2CL, an improvement of $4.6\%$, in Cl2Pr, an improvement of $1.4\%$ is observed by the addition of GMix augmentation.
@@
TABLE VIII: Accuracy comparison (%) in the Partial Domain Adaptation (PDA) scenario on the Office-Home dataset.
Methods U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU Fu et al. (2020) 50.9 74.2 78.4 62.2 64.1 72.5 63.5 47.9 78.3 72.4 54.7 78.9 66.5
DANCE Saito et al. (2020a) 53.6 73.2 84.9 70.8 67.3 82.6 70.0 50.9 84.8 77.0 55.9 81.8 71.1
DCC Li et al. (2021) 54.2 47.5 57.5 83.8 71.6 86.2 63.7 65.0 75.2 85.5 78.2 82.6 70.9
OVANet Saito and Saenko (2021a) 34.1 54.6 72.1 42.4 47.3 55.9 38.2 26.2 61.7 56.7 35.8 68.9 49.5
GATE Chen et al. (2022b) 55.8 75.9 85.3 73.6 70.2 83.0 72.1 59.5 84.7 79.6 63.9 83.8 74.0
@@
LEAD Qu et al. (2024) 58.2 83.1 87.0 70.5 75.4 83.3 73.7 50.4 83.7 78.3 58.7 83.2 73.8
+ GMix 58.9 76.3 87.7 71.3 76.8 83.4 73.9 59.2 81.9 79.1 63.3 80.6 74.3
TABLE IX: Top-1 (%) accuracy of self-supervised learning methods. Adding GMix yields better performance.
@@
MoCo v2 He et al. (2020) 80.31 40.82 51.36
+ GMix 82.36 42.27 53.57
SimSiam Chen and He (2021) 86.93 48.34 40.37
+ GMix 89.42 49.35 42.83
5.7 Self-Supervised Learning
@@
We experimented with self-supervised learning models MoCo v2 He et al. (2020) and SimSiam Chen and He (2021) by incorporating our proposed GMix augmentation method. Table IX showcases the Top-1 accuracy of self-supervised learning methods, comparing the performance of MoCo v2 and SimSiam, both with and without augmentation. On the Flower102, Stanford Cars, and Aircraft datasets, the baseline MoCo v2 achieves accuracies of $80.31\%$, $40.82\%$, and $51.36\%$, respectively. Incorporating GMix delivers absolute gains of $2.05\%$, $1.45\%$ and $2.21\%$ on these datasets. Similarly, SimSiam baseline achieved accuracies of $86.93\%$, $48.34\%$, and $40.37\%$ on the Flower102, Stanford Cars, and Aircraft datasets. Augmenting it with the proposed method gives absolute gains of $2.49\%$, $1.01\%$ and $2.46\%$. These results illustrate the efficacy of GMix, as it consistently boosts the performance across various datasets for self-supervised learning methods. The systematic gains highlight the robustness of our approach, demonstrating its potential to enhance the accuracy of state-of-the-art models in different domains.
6 Analysis and Discussion
@@
In this section, we present an in-depth analysis of the different design choices made in GMix, supporting it by a comprehensive ablation study.
6.1 Grad-CAM Visualization
@@
Gradient-weighted Class Activation Mapping (Grad-CAM) Selvaraju et al. (2017) is used to visualize class-discriminative regions in input images. It works by computing the gradients of the target class with respect to the feature maps in the last convolutional layer, highlighting the important regions contributing to the prediction. This helps in understanding where the model is focusing to make decisions, making it useful for interpretability. Using models trained from scratch, comparison with four existing methods Zhang et al. (2018); Yun et al. (2019b); Kim et al. (2020a); Kang and Kim (2023); Uddin et al. (2020) using Grad-CAM visualization is presented in Figure 4. We observe that the models with existing mixup based methods have less focused region in some images, compared to our proposed GMix. It shows that these methods focus specific object features such as front lights or wheels of the car only. The Grad-CAM saliency map of the model trained using GMix focuses on the full object, implying better generalization. This property can be attributed to the fact that GMix works with a single image and mixes different parts of the same image and its edited versions to augment the data. Our method covers more area of the object while focusing less on the context, whereas the compared approaches are more local within the object and in some cases also have focus outside the object. Thus GMix better retains the structure of the foreground object, resulting in an enlarged focus on the entire object representing better generalization.
@@
Figure 4: Representative Grad-CAM visualizations of saliency maps for Mixup Zhang et al. (2018), CutMix Yun et al. (2019b), PuzzleMix Kim et al. (2020a), GuidedMixup Kang and Kim (2023) methods on random input images of Stanford Cars dataset. We train ResNet50 from scratch for a fair comparison. Saliency maps for GMix are more consistent and better focused on the salient image regions.
6.2 Ablation Studies
@@
To assess the significance of the component of our approach, we performed an ablation study by systematically removing the impact of individual components and recording the performance on ResNet-50 across the Stanford Cars and Flowers-102 datasets, as shown in Table X.
@@
In the table, the baseline He et al. (2016b) using only the original images $I_{i}$ obtains Top-1 and Top-5 accuracies of $85.52\%$ $90.34\%$ on Stanford Cars and $78.83\%$ $94.38\%$ on Flowers102. Introducing fractal interpolation ($F_{v}$) to the input images brings a slight improvement in performance for both datasets. We then eliminate both seamless concatenation ($H_{iju}$) and fractal interpolation ($F_{v}$) by conducting experiments using only the edited images ($\hat{I}_{ij}$) as augmented images for training. This setup aligns more closely with methods in Trabucco et al. (2024); Azizi et al. (2023), which also use diffusion-generated images as augmentations. Our results for this experiment in the table are consistent with the insights of Azizi et al. (2023), suggesting that using generated images directly does not always significantly outperform the vanilla approach. We then perform experiments where edited images are interpolated with fractal images during training. In the Cars dataset, this leads to a slight improvement over the baseline, with Top-1 and Top-5 accuracies of $89.42\%$ and $91.57\%$, respectively. However, the Flowers dataset shows a slight drop in performance, indicating that fractal interpolation is more effective when combined with greater diversity in the training set. Next, we remove the fractal interpolation but retain the seamless concatenation of edited ($\hat{I}_{ij}$) and original ($I_{i}$) images to form hybrid images ($H_{iju}$) for data augmentation. This boosts performance for both datasets. We conjecture that this improvement is due to the combined presence of both generated and original image content in each augmented image, highlighting the importance of the seamless concatenation step in GMix. Finally, when all components, including fractal interpolation, are incorporated, the model achieves its best performance for both datasets. These consistent improvements with each additional component emphasize the effectiveness of the design choices in GMix for data augmentation.
@@
TABLE X: Ablation study using Stanford Cars (Cars) and Flowers102 (Flow) datasets. Top-1 and Top-5 accuracies are reported with different combinations of $I_{i}$: Input image, $\hat{I}_{ij}$: Edited images using prompts $p_{j}$, $H_{iju}$: Hybrid images using random mask $M_{u}$, and $F_{v}$: fractal images used to obtain final interpolated image $A_{ijuv}$.
$I_{i}$
$\hat{I}_{ij}$
$H_{iju}$
$F_{v}$
Cars Top-1 85.52 86.73 87.63 89.42 90.83 91.30
Top-5 90.34 92.38 90.23 91.57 97.56 99.96
Flow Top-1 78.73 78.34 77.38 77.81 79.31 80.27
Top-5 94.38 94.91 93.15 93.24 95.38 95.51
6.3 Augmentation Overhead
@@
We evaluate the computational overhead of various state-of-the-art image augmentation methods, including GMix, in relation to their performance improvements. As defined by Kang et al.Kang and Kim (2023), the augmentation overhead $\mathcal{A_{O}}$ is calculated as:
@@
where $\mathcal{T}_{aug}$ refers to the total time required for image generation and training, and $\mathcal{T}_{van}$ represents the training time of the baseline model without augmentation. Although GMix could leverage parallel processing to speed up image generation, in this experiment, we used sequential processing to ensure a fair comparison. As apparent in Fig. 5, GMix achieves an excellent balance between performance and augmentation overhead, surpassing all other methods in accuracy while maintaining significantly lower overhead compared to Co-Mixup Kim et al. (2020b) and SaliencyMix Uddin et al. (2020). Additionally, GMix can be further optimized by saving the generated images offline, enabling quicker execution of multiple training experiments, which is particularly useful for research and optimization purposes.
@@
6.4 Impact of Prompt Selection
Diffusion model outputs are dependent on the input prompts Du et al. (2023). Our approach to designing bespoke conditional prompts focuses on creating prompts that edit images while maintaining their structural integrity, making them applicable across various datasets. As outlined in the main text, we introduce filter-like prompts such as snowy, sunset, and rainbow, demonstrating their effectiveness in training robust classifiers. We also explore some bad prompts that are unsuitable for the image editing in GMix as shown in Figure 6. Descriptive or overly complex prompts tend to produce images that deviate significantly from the original distribution, often leading to unrealistic foregrounds and backgrounds, making them ineffective for training. It further emphasizes the value of our proposed filter-like bespoke conditional prompts, which avoid introducing unwanted alterations to the training data.
6.5 Replacing Concatenation with CutMix and Mixup
In these experiments, we integrate GMix with state-of-the-art image augmentation techniques Yun et al. (2019b); Zhang et al. (2018) at an appropriate component level to assess potential performance improvements. Specifically, we replace our seamless concatenation step with CutMix and Mixup, while keeping the rest of the GMix pipeline unchanged. For the CutMix + GMix approach, we substitute the concatenation step in GMix with the random cropping technique from CutMix. Here, a patch is randomly cropped from the generated image and pasted onto the original image. In the Mixup + GMix variant, we replace the concatenation with pixel-level blending of the original and generated images, as described in Zhang et al. (2018). The results are summarized in Table XI. Using either CutMix Yun et al. (2019b) or Mixup Zhang et al. (2018) yields improvements over the original techniques. While GMix provides consistent gains in these settings, the best performance of $81.30\%$ is achieved when our originally proposed method is used.
@@ 6.5 Replacing Concatenation with CutMix and Mixup
TABLE XI: Combining GMix with CutMix and Mixup image augmentation methods by replacing the seamless concatenation step of GMix with the masking techniques CutMix & Mixup.
Method Top-1 (%)
ResNet50505050 He et al. (2016b) 78.73
+ CutMix Yun et al. (2019b) 79.22
+ CutMix Yun et al. (2019b) + GMix 79.58
+ Mixup Zhang et al. (2018) 79.34
+ Mixup Zhang et al. (2018) + GMix 80.20
+ GMix 81.30
6.6 Increasing Smooth Mask Vs Performance
Experiments were also conducted using the Flowers102 dataset to assess how different smooth masking strategies affect the overall performance of the proposed method. The results are reported in Table XII. Employing even a single type of mask - specifically, vertical mask; our approach significantly outperforms the vanilla He et al. (2016b) baseline in terms of accuracy. Introducing both vertical and horizontal masks leads to further improvements. The highest accuracy is achieved when both mask types are combined with random flipping between the positions of input and generated images, enhancing the diversity of the training data. Moreover, the method is compatible with masking techniques from previous studies, such as those in Yun et al. (2019b); Zhang et al. (2018).
TABLE XII: Effects of masking in GMix on Flower102 dataset. All variants yield notably superior results compared to the vanilla ResNet-50. However, the best results are achieved when horizontal (Hor) and vertical (Ver) masks are used with flipping.
Mask Top-1 (%) Top-5 (%)
Vanilla He et al. (2016b) 89.74 94.38
Ver Mask ( ) 94.02 98.42
Hor + Ver Masks ( , ) 94.27 99.03
Hor + Ver + Flipping ( , , , ) 95.37 99.39
7 Conclusion
In this work, a data augmentation technique dubbed as GMix is proposed which is built on pre-trained image editing diffusion models. GMix enhances diversity in the training data while preserving the original semantic content of the input images. It employs editing, smooth concatenation, and fractal blending to produce the final augmented images. GMix has shown performance improvement across multiple tasks including general classification, fine-grained classification, domain adaptation, addressing data scarcity, fine-tuning, self-supervised learning, and adversarial robustness, over several architectures and benchmark datasets including ImageNet-1k, Tiny-ImageNet-200, CIFAR-100, Oxford Flower102, Caltech Birds, Stanford-Cars, FGVC Aircraft, and Office-Home. Across a wide range of experiments, our method has consistently demonstrated improved performance outperforming existing SOTA image augmentation methods. The proposed method has the potential to improve learning of deep models across a wide range of tasks.
Limitations: GMix has two key limitations: (1) It relies on text prompts for image editing, and inappropriate text prompts may result in undesired outputs. To address this, we propose a set of filter-like prompts that can be applied to a wide range of natural images. (2) GMix introduces additional overhead for editing images, but this cost may be considered as a trade off for improved convergence in large-scale learning models. Image editing cost can be effectively handled by generating edited images in advance and employing faster image editing methods.

### v4

@@ V-F Domain Adaptation
When DNNs encounter data from distributions that differ from their training data, a noticeable decline in performance is often observed. To tackle this issue, domain adaptation framework [93, 94, 95, 96, 97] has been developed, which utilizes labeled data from source domains to train DNNs for unlabeled target domains using a transductive learning approach. We validate the effectiveness of GenMix using LEAD framework [54] on OfficeHome [76] dataset. For a fair comparison, we access source-only and augment data with GenMix. We experiment with different settings including across partial-set domain adaptation (PDA), open-set domain adaptation (OSDA), and open-partial-set domain adaptation (OPDA). The OPDA results are reported in Table VII, while OSDA and PDA performance evaluations are summarized in Tables VIII and IX, respectively.
@@
Open-Partial Domain Adaptation (OPDA). OPDA [98, 99] falls under the broader scope of universal domain adaptation, where there is no prior knowledge about the label shift, such as common classes or the number of categories in the target domain. Some existing approaches [100, 101] require simultaneous access to both source and target data which may become impractical due to data protection regulations [102]. In Table VII, our baseline LEAD [54] has an average H-score of $75.0\%$, while LEAD + GenMix achieves $77.2\%$, representing a significant $2.2\%$ improvement in the average H-score. Particularly, for Ar2Pr, addition of our augmentation strategy improved performance of LEAD [54] by $4.9\%$, for CL2Ar the improvement is $6.2\%$ and for Re2Ar improvement is $4.4\%$. The results affirm that GenMix is an effective data augmentation for open partial domain adaptation scenarios.
@@
Open-set Domain Adaptation (OSDA). In OSDA a significant portion of the target samples comes from new categories that are absent in the source domain. Since these categories do not have specific labels, existing methods often treat them collectively as a single ”unknown” class. In Table VIII, LEAD + GenMix achieved an average improvement in the H-score of $1.5\%$ compared to the baseline LEAD, underscoring the significance of our proposed augmentation strategy. Particularly, considerable improvements of $4.1\%$ for Cl2Ar, $5.4\%$ for Re2CL and $3.4\%$ for Re2Pr; can be observed by the addition of GenMix augmentation in the baseline LEAD [54] method.
@@
Partial Domain Adaptation (PDA). PDA focuses on adapting a model trained on a source domain to a target domain where only a subset of the classes overlap. We compare the performance of a number of techniques on the Office-Home dataset while adapting LEAD as LEAD + GenMixto showcase the gain enabled by our method. Table IX shows the accuracy for various domain adaptation tasks as well as the average accuracy across all tasks. LEAD + GenMix achieved an average improvement in the H-score of $0.5\%$ over the LEAD. Particularly, for Pr2Cl maximum improvement of $9.5\%$, in Re2CL, an improvement of $4.6\%$, in Cl2Pr, an improvement of $1.4\%$ is observed by the addition of GenMix augmentation.
@@
TABLE IX: Accuracy comparison (%) in the Partial Domain Adaptation (PDA) scenario on the Office-Home dataset.
Method U SF CF Ar2Cl Ar2Pr Ar2Re Cl2Ar Cl2Pr Cl2Re Pr2Ar Pr2Cl Pr2Re Re2Ar Re2Cl Re2Pr Avg
CMU [86] 50.9 74.2 78.4 62.2 64.1 72.5 63.5 47.9 78.3 72.4 54.7 78.9 66.5
DANCE [87] 53.6 73.2 84.9 70.8 67.3 82.6 70.0 50.9 84.8 77.0 55.9 81.8 71.1
DCC [88] 54.2 47.5 57.5 83.8 71.6 86.2 63.7 65.0 75.2 85.5 78.2 82.6 70.9
OVANet [89] 34.1 54.6 72.1 42.4 47.3 55.9 38.2 26.2 61.7 56.7 35.8 68.9 49.5
GATE [90] 55.8 75.9 85.3 73.6 70.2 83.0 72.1 59.5 84.7 79.6 63.9 83.8 74.0
@@
LEAD [54] 58.2 83.1 87.0 70.5 75.4 83.3 73.7 50.4 83.7 78.3 58.7 83.2 73.8
+ GenMix 58.9 76.3 87.7 71.3 76.8 83.4 73.9 59.2 81.9 79.1 63.3 80.6 74.3
TABLE X: Top-1 (%) accuracy of self-supervised learning methods. Adding GenMix yields better performance.
@@
MoCo v2 [103] 80.31 40.82 51.36
DiffuseMix [53] 82.15 41.73 53.28
GenMix 83.57 43.71 54.68
SimSiam [104] 86.93 48.34 40.37
DiffuseMix [42] 89.24 49.17 42.63
GenMix 90.45 51.78 43.75
V-G Self-Supervised Learning
@@
We experimented with self-supervised learning models MoCo v2 [103] and SimSiam [104] by incorporating our proposed GenMix augmentation method. Table X showcases the Top-1 accuracy of self-supervised learning methods, comparing the performance of MoCo v2 and SimSiam, both with and without augmentation. On the Flower102, Stanford Cars, and Aircraft datasets, the baseline MoCo v2 achieves accuracies of GenMix delivers absolute gains of $1.42\%$, $1.98\%$ and $1.40\%$ on these datasets compared to DiffuseMix [53]. Similarly, SimSiam baseline achieved accuracies of $86.93\%$, $48.34\%$, and $40.37\%$ on the Flower102, Stanford Cars, and Aircraft datasets. Compared with DiffuseMix [53], augmenting it with the proposed method gives absolute gains of $1.21\%$, $2.61\%$ and $1.12\%$. These results illustrate the efficacy of GenMix, as it consistently boosts the performance across various datasets for self-supervised learning methods. The systematic gains highlight the robustness of our approach, demonstrating its potential to enhance the accuracy of state-of-the-art models in different domains.
VI Analysis and Discussion
@@
In this section, we present an in-depth analysis of the different design choices made in GenMix, supporting it by a comprehensive ablation study.
VI-A Grad-CAM Visualization
@@
Gradient-weighted Class Activation Mapping (Grad-CAM) [105] is used to visualize class-discriminative regions in input images. It works by computing the gradients of the target class with respect to the feature maps in the last convolutional layer, highlighting the important regions contributing to the prediction. This helps in understanding where the model is focusing to make decisions, making it useful for interpretability. Using models trained from scratch, comparison with four existing methods [28, 27, 19, 24, 22] using Grad-CAM visualization is presented in Figure 5. We observe that the models with existing mixup based methods have less focused region in some images, compared to our proposed GenMix. It shows that these methods focus specific object features such as front lights or wheels of the car only. The Grad-CAM saliency map of the model trained using GenMix focuses on the full object, implying better generalization. This property can be attributed to the fact that GenMix works with a single image and mixes different parts of the same image and its edited versions to augment the data. Our method covers more area of the object while focusing less on the context, whereas the compared approaches are more local within the object and in some cases also have focus outside the object. Thus GenMix better retains the structure of the foreground object, resulting in an enlarged focus on the entire object representing better generalization.
@@
Figure 5: Representative Grad-CAM visualizations of saliency maps for Mixup [28], CutMix [27], PuzzleMix [19], GuidedMixup [24] methods on random input images of Stanford Cars dataset. We train ResNet50 from scratch for a fair comparison. Saliency maps for GenMix are more consistent and better focused on the salient image regions.
VI-B Ablation Studies
@@
To assess the significance of the component of our approach, we performed an ablation study by systematically removing the impact of individual components and recording the performance on ResNet-50 across the Stanford Cars and Flowers-102 datasets, as shown in Table XI.
@@
In the table, the baseline [66] using only the original images $I_{i}$ obtains Top-1 and Top-5 accuracies of $85.52\%$ on Stanford Cars and $78.83\%$ on Flowers102. Introducing fractal interpolation ($F_{v}$) to the input images brings a slight improvement in performance for both datasets. We then eliminate both seamless concatenation ($H_{iju}$) and fractal interpolation ($F_{v}$) by conducting experiments using only the edited images ($\hat{I}_{ij}$) as augmented images for training. This setup aligns more closely with methods in [46, 47], which also use diffusion-generated images as augmentations. Our results for this experiment in the table are consistent with the insights of [47], suggesting that using generated images directly does not always significantly outperform the vanilla approach. We then perform experiments where edited images are interpolated with fractal images during training. In the Cars dataset, this leads to a slight improvement over the baseline, with Top-1 accuracies of $89.72\%$. However, the Flowers dataset shows a slight drop in performance, indicating that fractal interpolation is more effective when combined with greater diversity in the training set. Next, we remove the fractal interpolation but retain the seamless concatenation of edited ($\hat{I}_{ij}$) and original ($I_{i}$) images to form hybrid images ($H_{iju}$) for data augmentation. This boosts performance for both datasets. We conjecture that this improvement is due to the combined presence of both generated and original image content in each augmented image, highlighting the importance of the seamless concatenation step in GenMix. Finally, when all components, including fractal interpolation, are incorporated, the model achieves its best performance for both datasets. These consistent improvements with each additional component emphasize the effectiveness of the design choices in GenMix for data augmentation.
@@
TABLE XI: Ablation study using Stanford Cars (Cars) and Flowers102 (Flowers) datasets. Top-1 and Top-5 accuracies are reported with different combinations of $I_{i}$: Input image, $\hat{I}_{ij}$: Edited images using prompts $p_{j}$, $H_{iju}$: Hybrid images using random mask $M_{u}$, and $F_{v}$: fractal images used to obtain final interpolated image $A_{ijuv}$.
$I_{i}$ - - - -
$\hat{I}_{ij}$ - - - -
$H_{iju}$ - - - -
$F_{v}$ - - -
Cars
Top-1 85.52 86.73 88.13 89.72 91.87 92.48
Flowers
Top-1 78.73 78.34 77.88 77.81 90.67 81.56
VI-C Augmentation Overhead
@@
We evaluate the computational overhead of various state-of-the-art image augmentation methods, including GenMix, in relation to their performance improvements. As defined by Kang et al.[24], the augmentation overhead $\mathcal{A_{O}}$ is calculated as:
@@
where $\mathcal{T}_{aug}$ refers to the total time required for image generation and training, and $\mathcal{T}_{van}$ represents the training time of the baseline model without augmentation. Although GenMix could leverage parallel processing to speed up image generation, in this experiment, we used sequential processing to ensure a fair comparison. As apparent in Fig. 6, GenMix achieves an excellent balance between performance and augmentation overhead, surpassing all other methods in accuracy while maintaining significantly lower overhead compared to Co-Mixup [21] and SaliencyMix [22]. Additionally, GenMix can be further optimized by saving the generated images offline, enabling quicker execution of multiple training experiments, which is particularly useful for research and optimization purposes.
@@
VI-D Impact of Prompt Selection
Diffusion model outputs are dependent on the input prompts [41]. Our approach to designing bespoke conditional prompts focuses on creating prompts that edit images while maintaining their structural integrity, making them applicable across various datasets. As outlined in the main text, we introduce filter-like prompts such as snowy, sunset, and rainbow, demonstrating their effectiveness in training robust classifiers. We also explore some bad prompts that are unsuitable for the image editing in GenMix as shown in Figure 7. Descriptive or overly complex prompts tend to produce images that deviate significantly from the original distribution, often leading to unrealistic foregrounds and backgrounds, making them ineffective for training. It further emphasizes the value of our proposed filter-like bespoke conditional prompts, which avoid introducing unwanted alterations to the training data.
@@ VI-D Impact of Prompt Selection
VI-E Increasing Smooth Mask Vs Performance
Experiments were also conducted using the Flowers102 dataset to assess how different smooth masking strategies affect the overall performance of the proposed method. The results are reported in Table XII. Employing even a single type of mask - specifically, vertical mask; our approach significantly outperforms the vanilla [66] baseline in terms of accuracy. Introducing both vertical and horizontal masks leads to further improvements. The highest accuracy is achieved when both mask types are combined with random flipping between the positions of input and generated images, enhancing the diversity of the training data. Moreover, the method is compatible with masking techniques from previous studies, such as those in [27, 28].
TABLE XII: Effects of masking in GenMix on Flower102. All variants yield notably superior results compared to the vanilla. However, best results are achieved when horizontal (Hor) and vertical (Ver) and patchswap masks are used with vice versa.
Mask Top-1 (%)
Vanilla [66] 89.74
Ver Mask ( ) 94.11
Hor + Ver Masks ( , ) 94.23
Hor + Ver + Flipping ( , , , ) 95.55
Hor + Ver + Flipping + PatchSwap ( , ) 96.26
VII Conclusion
In this work, a data augmentation technique dubbed as GenMix is proposed which is built on pre-trained image editing diffusion models. GenMix enhances diversity in the training data while preserving the original semantic content of the input images. It employs editing, smooth concatenation, and fractal blending to produce the final augmented images. GenMix has shown performance improvement across multiple tasks including general classification, fine-grained classification, domain adaptation, addressing data scarcity, fine-tuning, self-supervised learning, and adversarial robustness, over several architectures and benchmark datasets including ImageNet-1k, Tiny-ImageNet-200, CIFAR-100, Oxford Flower102, Caltech Birds, Stanford-Cars, FGVC Aircraft, and Office-Home. Across a wide range of experiments, our method has consistently demonstrated improved performance outperforming existing SOTA image augmentation methods. The proposed method has the potential to improve learning of deep models across a wide range of tasks.
Limitations: GenMix has two key limitations: (1) It relies on text prompts for image editing, and inappropriate text prompts may result in undesired outputs. To address this, we propose a set of filter-like prompts that can be applied to a wide range of natural images. (2) GenMix introduces additional overhead for editing images, but this cost may be considered as a trade off for improved convergence in large-scale learning models. Image editing cost can be effectively handled by generating edited images in advance and employing faster image editing methods.